# Production of high-spin $\omega_{J}/\rho_{J}$ ($J=3,4,5$) mesons in $\pi^{-}p$ reactions

### v1

@@ Abstract
In this work, we perform a comprehensive investigation of the production of high‑spin $\omega_{J}$ and $\rho_{J}$ mesons ($J=3,4,5$) in $\pi^{-}p$ reactions using an effective Lagrangian approach. By constructing the relevant tt‑channel processes and calibrating the model with a single adjustable parameter fitted to existing data, we successfully reproduce the measured total and differential cross sections for the $J=3$ states $\omega_{3}(1670)$ and $\rho_{3}(1690)$. Within the same framework, we predict the production cross sections for the higher‑spin partners $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, and $\rho_{5}(2350)$. Our results show that these states exhibit measurable cross sections with characteristically forward‑peaked angular distributions, underscoring their strong potential for observation in future $\pi p$ meson‑beam experiments.
@@ I INTRODUCTION
A particularly interesting subset of the light-hadron family comprises mesons with high spin, such as the $\rho_{J}$ and $\omega_{J}$ states ($J=3,4,5$). Their systematics can be summarized, as shown in Fig. 1. However, precisely because of their high spin, detailed theoretical studies of these states remain relatively scarce Anisovich et al. (2000); He et al. (2013); Bai et al. (2025); Li et al. (2023); Wang et al. (2022b); Pang et al. (2015). Their production dynamics are thus poorly understood, making them a pertinent target for investigation.
@@
Figure 1: A concise summary of light mesons.
In this work, we investigate the production of high-spin $\omega_{J}$ and $\rho_{J}$ mesons using an effective Lagrangian approach combined with Reggeization. The study proceeds in two stages. First, we constrain our model parameters using available data for the production of the better-established $J=3$ states, $\omega_{3}(1670)$ and $\rho_{3}(1690)$, in $\pi^{-}p$ reactions. The sole free parameter is a universal cutoff, $\Lambda_{t}$, in the form factor. Second, employing the same model and the determined $\Lambda_{t}$, we present predictions for the total and differential cross sections of their higher-spin partners: $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, and $\rho_{5}(2350)$, whose production has not yet been measured. This approach provides a consistent, nearly parameter-free link between the known $J=3$ sector and the unexplored $J=4,5$ sector, offering concrete guidance for future experimental searches.
The paper is organized as follows. Following this Introduction, Section II focuses on the production of $\omega_{3}(1670)$ and $\rho_{3}(1690)$ in $\pi p$ scattering, where the model is shown to describe the available experimental data well. Based on these results, Section III presents predictions for the production characteristics of the higher-spin $\omega$ and $\rho$ states with $J=4$ and $J=5$. Finally, Section IV summarizes the work and its implications.
@@ II Production of $\omega_{3}(1670)$ and $\rho_{3}(1690)$ in $\pi p$ reactions
The Regge trajectory model has been successfully applied to describe hadron production in the high-energy region Ozaki et al. (2010); Storrow (1984). When employing conventional Feynman propagators, the resulting cross sections show a monotonically increasing dependence on the center-of-mass energy, which is inconsistent with the corresponding experimental measurements shown in Figs. 3 and 5. To address this problem, in the present framework, Reggeization is implemented by replacing the $t$-channel Feynman propagators in the Feynman amplitudes Eqs. (7) and (8) with the corresponding Regge propagators:
@@
where the scale parameter $s_{\text{scale}}$ is fixed at 1 GeV. The Regge trajectories for the $\rho$ and $\pi$ are taken from Guidal et al. (1997); Wang et al. (2025):
$\begin{split}\alpha_{\rho}(t)=1+0.8(t-m_{\rho}^{2}),\end{split}$ (12)
$\begin{split}\alpha_{\pi}(t)=0.7(t-m_{\pi}^{2}).\end{split}$ (13)
@@
At $E_{\mathrm{cm}}=4.84\ \mathrm{GeV}$, the theoretical prediction (orange solid curve) agrees excellently with the data over the full angular range, accurately reproducing a characteristic dip near $\cos\theta\approx 0.98$. At $E_{\mathrm{cm}}=5.39\ \mathrm{GeV}$, slight deviations appear at very forward angles (around $\cos\theta\approx 0.97$). Nevertheless, the overall shape of the angular distribution, particularly its pronounced forward peaking, is well reproduced by the calculation. Although not all data points are encompassed at this higher energy, the theoretical curve captures the essential features of the measured distribution.
@@
A more detailed angular analysis is provided in Fig. 6, which shows the differential cross section $d\sigma/d\cos\theta$ at $\sqrt{s}=3.49\ \mathrm{GeV}$. The theoretical prediction (solid orange curve) and its associated uncertainty band ( corresponding to $\Lambda_{t}=3.5\pm 0.5\ \mathrm{GeV}$) are compared with the experimental data from Ref. Engler et al. (1974). The predicted angular distribution exhibits a pronounced forward peak ($\cos\theta\to 1$). Excellent agreement with the data is achieved in this forward region, as clearly shown in the magnified inset.
@@ III Production of higher-spin $\omega_{J}/\rho_{J}$ ($J=4,5$) mesons in $\pi^{-}p$ reactions
We now extend the same theoretical framework to predict the production cross sections of the higher-spin $\omega_{J}$ and $\rho_{J}$ mesons ($J=4,5$) in $\pi^{-}p$ reactions. This section is organized as follows. After a brief review of the experimental status of these states, we outline the theoretical formalism, including the effective Lagrangians and production amplitudes. Finally, we present the predicted total and differential cross sections.
The resonance parameters for the $\omega_{J}$ and $\rho_{J}$ mesons with $J=4$ and $5$ are summarized in Table 3. These states are the natural higher-spin partners of the $J=3$ $\omega_{3}(1670)$ and $\rho_{3}(1690)$. According to the PDG Navas et al. (2024), the $\omega_{4}(2250)$, $\rho_{4}(2230)$, and $\omega_{5}(2250)$ are categorized as “further states,” requiring more confirmation, whereas the $\rho_{5}(2350)$ is more firmly established.
Experimentally, these states have been observed in specific production channels. The $\omega_{4}(2250)$ was seen in $p\bar{p}$ annihilation, e.g., $p\bar{p}\to\omega\eta$ and $p\bar{p}\to\omega\pi^{0}\pi^{0}$ Anisovich et al. (2002a). Similarly, the $\rho_{4}(2230)$ was found in $\bar{p}\to\omega\pi^{0}, pp$$\bar{p}\to\omega\eta\pi^{0}$, and $p\bar{p}\to\pi^{+}\pi^{-}$ Anisovich et al. (2002b). The $\omega_{5}(2250)$ was only seen in $p\bar{p}\to\omega\eta$ and $p\bar{p}\to\omega\pi^{0}\pi^{0}$ Anisovich et al. (2001, 2002a), while the $\rho_{5}(2350)$ has been identified in several reactions, including $\pi^{-}p\to\omega\pi^{0}n$, $p\bar{p}\to\pi\pi$, and $p\bar{p}\to K^{+}K^{-}$ Alde et al. (1994); Hasan and Bugg (1994); Carter (1978).
Several features can be inferred from the PDG values listed in Table 3. The isoscalar $\omega_{4}(2250)$ and the isovector $\rho_{4}(2230)$ have comparable masses around 2.24GeV2.24\ $\mathrm${GeV}, with the $\rho_{4}$ being slightly lighter. Their total widths are of order 150150210MeV210\ $\mathrm${MeV}. For $J=5$, the $\omega_{5}(2250)$ and $\rho_{5}(2350)$ are heavier and exhibit significantly larger widths. Notably, the mass splitting between $\omega_{J}$ and $\rho_{J}$ partners is small for $J=4$, but becomes more pronounced for $J=5$, reaching about 8080100MeV100\ $\mathrm${MeV}, with the $\rho_{5}$ being the heavier state.
Table 3: The experimental information of the $\omega_{J}/\rho_{J}$ ($J=4,5$) mesons. Here, the masses and widths are average values taken from PDG Navas et al. (2024).
@@
We now examine the theoretical decay properties of the higher-spin $\omega_{J}$ and $\rho_{J}$ ($J=4,5$) mesons. Their predicted partial decay channels and corresponding branching ratios, calculated within the QPC model Pang et al. (2015), are compiled in Table 4.
@@
Production mechanism. Building on the successful description of $\omega_{3}(1670)$ and $\rho_{3}(1690)$ production in $\pi^{-}p$ scattering using a universal cutoff $\Lambda_{t}=3.5\pm 0.5\ \mathrm{GeV}$, the similiar framework is further employed to predict the production of the corresponding higher-spin partners.
The phenomenological basis for this extension lies in the decay patterns predicted by the QPC model (Table 4). Although not always the dominant modes, decays such as $\omega_{4}\to\pi\rho$ and $\rho_{4}\to\pi a_{1}$ occur with non-negligible branching ratios. This indicates that these higher-spin states can be produced in $\pi^{-}p$ reactions via $t$-channel single-meson exchange, in direct analogy to the $J=3$ case. The choice of the exchanged meson in the $t$-channel is guided by the dominant $\pi$-containing decay modes predicted by the QPC model (Table 4). The corresponding tree-level Feynman diagrams for the reactions $\pi^{-}p\to\omega_{4}(2250)n$, $\pi^{-}p\to\rho_{4}(2230)n$, $\pi^{-}p\to\omega_{5}(2250)n$, and $\pi^{-}p\to\rho_{5}(2350)^{-}p$ are depicted in Figs. 7(a)–(d), respectively. Given the limited existing experimental data on their production, these parameter-free predictions offer valuable guidance for future investigations in this energy region.
Table 4: The partial decay channels of the $\omega_{J}/\rho_{J}$ ($J=4,5$) mesons Pang et al. (2015).
$\omega_{4}(2250)$ $\rho_{4}(2230)$ $\omega_{5}(2250)$ $\rho_{5}(2350)$
Channel BR Channel BR Channel BR Channel BR
$a_{1}\rho$ 0.35 $\rho b_{1}$ 0.27 $\rho a_{2}$ 0.58 $\rho f_{2}$ 0.24
$a_{2}\rho$ 0.30 $\rho\rho$ 0.20 $\omega f_{2}$ 0.23 $\omega a_{2}$ 0.24
$\omega f_{2}$ 0.13 $\rho f_{2}$ 0.10 $\pi b_{1}$ 0.09 $\rho b_{1}$ 0.13
$\pi\rho$ 0.06 $\pi a_{1}$ 0.09 $\pi a_{1}$ 0.04 $\rho\rho$ 0.09
$\pi\rho(1700)$ 0.05 $\omega a_{1}$ 0.09 $\rho\pi$ 0.03 $\pi\pi_{2}$ 0.07
$\omega f_{1}$ 0.04 $\omega a_{2}$ 0.06 $\eta h_{1}$ 0.01 $\pi a_{2}$ 0.07
$\rho\rho$ 0.03 $\pi a_{2}$ 0.04 $\pi h_{1}$ 0.05
$\pi\omega$ 0.01 $\pi a_{1}$ 0.04
$\pi\omega$ 0.01
@@
[Fig. 7(a)]. This reaction proceeds via $\rho$-meson exchange, motivated by the fact that $\pi\rho$ is the most prominent $\pi$-containing decay of $\omega_{4}(2250)$ (Table 4). The effective Lagrangian for the $W_{4}V_{1}P$ vertex is
@@
[Fig. 7(b)]. The dominant $\pi$-containing decay $\pi a_{1}$ leads to $a_{1}$-meson exchange. The effective Lagrangian is
@@
[Fig. 7(c)]. This channel is dominated by $t$-channel $b_{1}$ exchange. The corresponding Lagrangian is
@@
$\pi^{-}p\to\rho_{5}(2350)n$
[Fig. 7(d)]. Here the exchange of an $h_{1}$ meson is involved, following the QPC prediction that $\pi h_{1}$ is the most important $\pi$-containing channel. The effective Lagrangian is
@@
$g_{\omega_{4}\rho\pi}=3.91\ \mathrm{GeV}^{-2}$ $g_{\rho NN}=2.05$
$f_{\rho NN}=13.40$
@@
Using the effective Lagrangians introduced in the previous section, we write the tree-level amplitudes for the four reactions as follows.
@@
Figure 8: The energy dependence of the total cross section for production of the $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, $\rho_{5}(2350)$ through $t$-channel with cutoff $\Lambda_{t}=3.5\pm 0.5$ GeV.

### v2

# Production of high-spin $\omega_{J}/\rho_{J}$ ($J=2,3,4,5$) mesons in $\pi^{-}p$ reactions

@@ Abstract
In this work, we perform a comprehensive investigation of the production of high‑spin $\omega_{J}$ and $\rho_{J}$ mesons ($J=2,3,4,5$) in $\pi^{-}p$ reactions using an effective Lagrangian approach. By constructing the relevant tt‑channel processes and calibrating the model with a single adjustable parameter fitted to existing data, we successfully reproduce the measured total and differential cross sections for the $J=3$ states $\omega_{3}(1670)$ and $\rho_{3}(1690)$. Within the same framework, we predict the production cross sections for their lower- and higher-spin partners: $\omega_{2}(1975)$, $\rho_{2}(1940)$, $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2350)$, and $\rho_{5}(2350)$. Our results show that these states exhibit measurable cross sections with characteristically forward‑peaked angular distributions, underscoring their strong potential for observation in future $\pi p$ meson‑beam experiments.
@@ I INTRODUCTION
A particularly interesting subset of the light-hadron family is that of high-spin mesons, such as the $\rho_{J}$ and $\omega_{J}$ states with $J=2,3,4,5$. Their systematics can be summarized, as shown in Fig. 1. However, precisely because of their high spin, detailed theoretical studies of these states remain relatively scarce Anisovich et al. (2000); He et al. (2013); Bai et al. (2025); Li et al. (2023); Wang et al. (2022b); Pang et al. (2015). Their production dynamics are thus poorly understood, making them a pertinent target for investigation.
@@
Figure 1: A concise summary of light mesons Navas et al. (2024).
In this work, we investigate the production of high-spin $\omega_{J}$ and $\rho_{J}$ mesons using an effective Lagrangian approach combined with Reggeization. The study proceeds in two stages. First, we constrain our model parameters using available data for the production of the better-established $J=3$ states, $\omega_{3}(1670)$ and $\rho_{3}(1690)$, in $\pi^{-}p$ reactions. The sole free parameter is a universal cutoff, $\Lambda_{t}$, in the form factor. Second, employing the same model and the extracted value of $\Lambda_{t}$, we present predictions for the total and differential cross sections of their lower- and higher-spin partners: $\omega_{2}(1975)$, $\rho_{2}(1940)$, $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2350)$, and $\rho_{5}(2350)$, whose production has not yet been measured. This approach provides a consistent, nearly parameter‑free link between the known $J=3$ sector and the unexplored $J=2,4,5$ sectors, offering concrete guidance for future experimental searches.
The paper is organized as follows. Following this Introduction, Section II focuses on the production of $\omega_{3}(1670)$ and $\rho_{3}(1690)$ in $\pi p$ scattering, where the model is shown to describe the available experimental data well. Based on these results, Section III presents predictions for the production characteristics of the $\omega$ and $\rho$ states with $J=2,4,5$. Finally, Section IV summarizes the work and its implications.
@@ II Production of $\omega_{3}(1670)$ and $\rho_{3}(1690)$ in $\pi p$ reactions
The Regge trajectory model has been successfully applied to describe hadron production in the high-energy region Ozaki et al. (2010); Storrow (1984). The motivation for replacing the usual Feynman propagator with a Regge propagator is to effectively account for the exchange of high-spin and high-mass states in the $t$ channel, whose contributions cannot be neglected at higher energies. When employing conventional Feynman propagators, the resulting cross sections show a monotonically increasing dependence on the center-of-mass energy, which is inconsistent with the corresponding experimental measurements shown in Figs. 3 and 5. To address this problem, in the present framework, Reggeization is implemented by replacing the $t$-channel Feynman propagators in the Feynman amplitudes Eqs. (7) and (8) with the corresponding Regge propagators Guidal et al. (1997):
@@
where the scale parameter $s_{\text{scale}}$ is fixed at 1 GeV, The functions $\alpha_{\rho}(t)$ and $\alpha_{\pi}(t)$ denote the $(J,M^{2})$ Regge trajectories Guidal et al. (1997); Wang et al. (2025):
$\begin{split}\alpha_{\rho}(t)=1+\alpha_{\rho}^{\prime}(t-m_{\rho}^{2}),\end{split}$ (12)
$\begin{split}\alpha_{\pi}(t)=\alpha_{\pi}^{\prime}(t-m_{\pi}^{2}).\end{split}$ (13)
where $\alpha_{\rho}^{\prime}=0.8\,\text{GeV}^{-2}$ and $\alpha_{\pi}^{\prime}=0.7\,\text{GeV}^{-2}$ are the slopes of the corresponding Regge trajectories, which can be generally constrained by the experimentally established states lying on the same trajectories Guidal et al. (1997); Wang et al. (2025).
@@
At $E_{\mathrm{cm}}=4.84\ \mathrm{GeV}$, the theoretical prediction (green solid curve) agrees excellently with the data over the full angular range, accurately reproducing a characteristic dip near $\cos\theta\approx 0.98$. At $E_{\mathrm{cm}}=5.39\ \mathrm{GeV}$, slight deviations appear at very forward angles (around $\cos\theta\approx 0.97$). Nevertheless, the overall shape of the angular distribution, particularly its pronounced forward peaking, is well reproduced by the calculation. Although not all data points are encompassed at this higher energy, the theoretical curve captures the essential features of the measured distribution.
@@
A more detailed angular analysis is provided in Fig. 6, which shows the differential cross section $d\sigma/d\cos\theta$ at $\sqrt{s}=3.49\ \mathrm{GeV}$. The theoretical prediction (solid green curve) and its associated uncertainty band ( corresponding to $\Lambda_{t}=3.5\pm 0.5\ \mathrm{GeV}$) are compared with the experimental data from Ref. Engler et al. (1974). The predicted angular distribution exhibits a pronounced forward peak ($\cos\theta\to 1$). Excellent agreement with the data is achieved in this forward region, as clearly shown in the magnified inset.
@@
We note that the structures in the effective Lagrangians for high-spin fields are not uniquely determined. In this work, we adopt the minimal derivative rule: for each vertex, we retain the lowest number of derivatives consistent with Lorentz invariance and parity conservation. The coupling constants are fixed by the corresponding partial decay widths. Hence, any alternative Lagrangian form that reproduces the same width would yield a proportionally adjusted coupling, leaving the overall scale of the cross section largely unchanged. The validity of our minimal choice at low energies is explicitly verified by the successful fit to the $J=3$ total and differential cross section data.
At high energies, the Reggeization of $t$-channel propagators further reduces the sensitivity to vertex details. The Regge amplitude is dominated by the universal Regge trajectory $\alpha(t)$, which governs the energy dependence, while the vertex structure contributes only to the residue. Consequently, the predicted production features for the $J=4,5$ states are robust against moderate variations in the Lagrangian construction, as shown in the following section.
@@ III Production of higher-spin $\omega_{J}/\rho_{J}$ ($J=4,5$) mesons in $\pi^{-}p$ reactions
We now extend the same theoretical framework to predict the production cross sections of the lower-spin and higher-spin $\omega_{J}$ and $\rho_{J}$ mesons ($J=2,4,5$) in $\pi^{-}p$ reactions. This section is organized as follows. After a brief review of the experimental status of these states, we outline the theoretical formalism, including the effective Lagrangians and production amplitudes. Finally, we present the predicted total and differential cross sections.
The resonance parameters for the $\omega_{J}$ and $\rho_{J}$ mesons with $J=2,3,4$ and $5$ are summarized in Table 3. These states are the natural lower-spin and higher-spin partners of the $J=3$ $\omega_{3}(1670)$ and $\rho_{3}(1690)$. According to the PDG Navas et al. (2024), the $\omega_{2}(1975)$, $\rho_{2}(1940)$, $\omega_{4}(2250)$, $\rho_{4}(2230)$, and $\omega_{5}(2250)$ are categorized as “further states,” requiring more confirmation, whereas the $\rho_{5}(2350)$ is more firmly established.
Experimentally, these states have been observed in specific production channels. The $\omega_{2}(1975)$ and $\omega_{4}(2250)$ are seen in $p\bar{p}$ annihilation, e.g., $p\bar{p}\to\omega\eta$ and $p\bar{p}\to\omega\pi^{0}\pi^{0}$ Anisovich et al. (2002a). Similarly, the $\rho_{2}(1940)$ and $\rho_{4}(2230)$ are found in $\bar{p}\to\omega\pi^{0}, pp$$\bar{p}\to\omega\eta\pi^{0}$, and $p\bar{p}\to\pi^{+}\pi^{-}$ Anisovich et al. (2002b). The $\omega_{5}(2250)$ was only seen in $p\bar{p}\to\omega\eta$ and $p\bar{p}\to\omega\pi^{0}\pi^{0}$ Anisovich et al. (2001, 2002a), while the $\rho_{5}(2350)$ has been identified in several reactions, including $\pi^{-}p\to\omega\pi^{0}n$, $p\bar{p}\to\pi\pi$, and $p\bar{p}\to K^{+}K^{-}$ Alde et al. (1994); Hasan and Bugg (1994); Carter (1978).
Several features can be inferred from the PDG values listed in Table 3. For $J=2$, the isoscalar $\omega_{2}(1975)$ has a mass of $1975\pm 20$ MeV and a width of $175\pm 25$ MeV, while the isovector $\rho_{2}(1940)$ has a mass of $1940\pm 40$ MeV and a width of $155\pm 40$ MeV. The mass splitting between $\omega_{2}$ and $\rho_{2}$ is approximately $35$ MeV, which is relatively small, and their widths are comparable. Moving to $J=4$, the isoscalar $\omega_{4}(2250)$ and the isovector $\rho_{4}(2230)$ have comparable masses around 2.24GeV2.24\ $\mathrm${GeV}, with the $\rho_{4}$ being slightly lighter. Their total widths are of order 150150210MeV210\ $\mathrm${MeV}. For $J=5$, the $\omega_{5}(2250)$ and $\rho_{5}(2350)$ are heavier and exhibit significantly larger widths. Notably, the mass splitting between $\omega_{J}$ and $\rho_{J}$ partners is small for $J=4$, but becomes more pronounced for $J=5$, reaching about 80MeV80\ $\mathrm${MeV}, with the $\rho_{5}$ being the heavier state.
Table 3: The experimental information of the $\omega_{J}/\rho_{J}$ ($J=2,4,5$) mesons. Here, the masses and widths are average values taken from the PDG Navas et al. (2024).
@@
We now examine the theoretical decay properties of the $\omega_{J}$ and $\rho_{J}$ ($J=2,4,5$) mesons. Their predicted partial decay channels and corresponding branching ratios, calculated within the QPC model Pang et al. (2015), are compiled in Table 4.
For the $\omega_{2}(1975)$, the dominant decay channel is $\pi\rho(1450)$ with a branching ratio of $0.48$, indicating a strong preference for decay into a pseudoscalar meson and an excited vector meson. The channel $\pi\rho$ also contributes significantly with a branching ratio of $0.24$. Other notable channels include $\pi b_{1}$ ($0.11$) and $\pi\rho_{3}(1690)$ ($0.07$). For the $\rho_{2}(1940)$, the largest decay channel is $\pi a_{2}(1320)$ with a branching ratio of $0.25$, highlighting decays into a pseudoscalar and a tensor meson. The channel $\pi\omega(1420)$ has a branching ratio of $0.21$, and the $\pi\omega$ contributes $0.15$. Other significant channels include $\pi h_{1}(1170)$ and $\rho\eta$.
@@
Production mechanism. Building on the successful description of $\omega_{3}(1670)$ and $\rho_{3}(1690)$ production in $\pi^{-}p$ scattering using a universal cutoff $\Lambda_{t}=3.5\pm 0.5\ \mathrm{GeV}$, the similiar framework is further employed to predict the production of the corresponding lower-spin and higher-spin partners.
The phenomenological basis for this extension lies in the decay patterns predicted by the QPC model (Table 4). Although not always the dominant modes, decays such as $\omega_{4}\to\pi\rho$ and $\rho_{4}\to\pi a_{1}$ occur with non-negligible branching ratios. This indicates that these states can be produced in $\pi^{-}p$ reactions via $t$-channel single-meson exchange, in direct analogy to the $J=3$ case. The choice of the exchanged meson in the $t$-channel is guided by the dominant $\pi$-containing decay modes predicted by the QPC model (Table 4). The corresponding tree-level Feynman diagrams for the reactions $\pi^{-}p\to\omega_{2}(1975)n$, $\pi^{-}p\to\rho_{2}(1940)p$, $\pi^{-}p\to\omega_{4}(2250)n$, $\pi^{-}p\to\rho_{4}(2230)n$, $\pi^{-}p\to\omega_{5}(2250)n$, and $\pi^{-}p\to\rho_{5}(2350)^{-}p$ are depicted in Fig. 7(a)-(f), respectively. Given the limited existing experimental data on their production, these parameter-free predictions offer valuable guidance for future investigations in this energy region.
Table 4: The partial decay channels of the $\omega_{J}/\rho_{J}$ ($J=2,4,5$) mesons Pang et al. (2015).
$\omega_{2}(1975)$ $\rho_{2}(1940)$ $\omega_{4}(2250)$ $\rho_{4}(2230)$ $\omega_{5}(2250)$ $\rho_{5}(2350)$
Channel BR Channel BR Channel BR Channel BR Channel BR Channel BR
$\pi\rho(1450)$ 0.48 $\pi a_{2}(1320)$ 0.25 $a_{1}\rho$ 0.35 $\rho b_{1}$ 0.27 $\rho a_{2}$ 0.58 $\rho f_{2}$ 0.24
$\pi\rho$ 0.24 $\pi\omega(1420)$ 0.21 $a_{2}\rho$ 0.30 $\rho\rho$ 0.20 $\omega f_{2}$ 0.23 $\omega a_{2}$ 0.24
$\pi b_{1}$ 0.11 $\pi\omega$ 0.15 $\omega f_{2}$ 0.13 $\rho f_{2}$ 0.10 $\pi b_{1}$ 0.09 $\rho b_{1}$ 0.13
$\pi\rho_{3}(1690)$ 0.07 $\pi h_{1}(1170)$ 0.10 $\pi\rho$ 0.06 $\pi a_{1}$ 0.09 $\pi a_{1}$ 0.04 $\rho\rho$ 0.09
$\omega\eta$ 0.03 $\rho\eta$ 0.07 $\pi\rho(1700)$ 0.05 $\omega a_{1}$ 0.09 $\rho\pi$ 0.03 $\pi\pi_{2}$ 0.07
$\pi\rho(1700)$ 0.02 $\pi a_{0}(1450)$ 0.06 $\omega f_{1}$ 0.04 $\omega a_{2}$ 0.06 $\eta h_{1}$ 0.01 $\pi a_{2}$ 0.07
$KK^{*}$ 0.02 $\rho\rho$ 0.05 $\rho\rho$ 0.03 $\pi a_{2}$ 0.04 $\pi h_{1}$ 0.05
$KK^{*}(1410)$ 0.02 $KK^{*}$ 0.04 $\pi\omega$ 0.01 $\pi a_{1}$ 0.04
$KK^{*}_{2}(1430)$ 0.03 $\pi\omega$ 0.01
$\rho\eta^{\prime}$ 0.03
@@
$\pi^{-}p\to\omega_{2}(1975)n$
[Fig. 7(a)]. This reaction proceeds via $\rho$-meson exchange, motivated by the fact that $\pi\rho$ is the most prominent $\pi$-containing decay of $\omega_{2}(1975)$ (Table 4). The effective Lagrangian for the $W_{2}V_{1}P$ vertex is
$\mathcal{L}_{W_{2}V_{1}P}=g_{W_{2}V_{1}P}\operatorname{tr}\left[W_{2}^{\alpha\beta}\left\{V_{\alpha},\partial_{\beta}P\right\}_{+}\right],$ (15)
where the fields $W_{2}$, $V_{1}$, and $P$ correspond to the spin-$2$ meson with $J^{PC}=2^{--}$, vector meson and pseudoscalar meson, respectively. The nonet of $2^{--}$ reads
$W_{2}^{\mu\nu}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{ccc}\frac{\omega_{2}^{\mu\nu}+\rho_{2}^{0\mu\nu}}{\sqrt{2}}&\rho_{2}^{+\mu\nu}&K_{2}^{+\mu\nu}\\ \rho_{2}^{-\mu\nu}&\frac{\omega_{2}^{\mu\nu}-\rho_{2}^{0\mu\nu}}{\sqrt{2}}&K_{2}^{0\mu\nu}\\ K_{2}^{-\mu\nu}&\bar{K}_{2}^{0\mu\nu}&\omega_{2}^{\mu\nu}\end{array}\right).$
The specific form of the $\omega_{2}\rho\pi$ vertex is as follows:
$\mathcal{L}_{\omega_{2}\rho\pi}=\frac{1}{2}g_{\omega_{2}\rho\pi}\omega_{2}^{\alpha\beta}\rho_{\alpha}^{+}\partial_{\beta}\pi^{-}.$ (16)
The $\rho NN$ vertex is presented in Eq. (3).
$\pi^{-}p\to\rho_{2}(1940)p$
[Fig. 7(b)]. The forms of the Lagrangian for the two vertices in this process are exactly the same as that in the previous process ($\pi^{-}p\to\omega_{2}(1975)n$), with only the exchanged particle changed from $\rho$ to $\omega$. Hence, we omit the Lagrangian required for this Feynman process here.
@@
[Fig. 7(c)]. This reaction proceeds via $\rho$-meson exchange, motivated by the fact that $\pi\rho$ is the most prominent $\pi$-containing decay of $\omega_{4}(2250)$ (Table 4). The effective Lagrangian for the $W_{4}V_{1}P$ vertex is
@@
[Fig. 7(d)]. The dominant $\pi$-containing decay $\pi a_{1}$ leads to $a_{1}$-meson exchange. The effective Lagrangian is
@@
[Fig. 7(e)]. This channel is dominated by $t$-channel $b_{1}$ exchange. The corresponding Lagrangian is
@@
$\pi^{-}p\to\rho_{5}(2350)p$
[Fig. 7(f)]. Here the exchange of an $h_{1}$ meson is involved, following the QPC prediction that $\pi h_{1}$ is the most important $\pi$-containing channel. The effective Lagrangian is
@@
$g_{\rho_{2}\omega\pi}=5.89$ $g_{\omega NN}=11.04$
$f_{\omega NN}=-2.02$
$g_{\omega_{2}\rho\pi}=4.39$ $g_{\rho NN}=2.05$
$g_{\omega_{4}\rho\pi}=3.91\ \mathrm{GeV}^{-2}$ $f_{\rho NN}=13.40$
@@
Using the effective Lagrangians introduced in the previous section, we write the tree-level amplitudes for the six reactions as follows.
$\pi^{-}p\to\omega_{2}(1975)n$ via $t$-channel $\rho$ exchange:
[…]
$\pi^{-}p\to\rho_{2}(1940)p$ via $t$-channel $\omega$ exchange:
[…]
@@
Figure 8: The energy dependence of the total cross section for production of the $\omega_{2}(1975)$, $\rho_{2}(1940)$, $\omega_{4}(2250)$, $\rho_{4}(2230)$, $\omega_{5}(2250)$, $\rho_{5}(2350)$ through $t$-channel with cutoff $\Lambda_{t}=3.5\pm 0.5$ GeV.